What is the enthalpy change for 3 mol liquid benzene transformed to vapour state at boiling temperature 800C and 1 bar pressure?
 

Grade:11

1 Answers

Naveen Kumar
9 years ago
change in enthalpy, dH = dU + pdV= dU + dn*R*T
As the change is occuring at the boiling temperature, dH = 3*R*T
use T = 273+80=353K.